OpenStudy (cookiemonster2412):

Find the equation of the line given the following information. two points: (-3,-1) (0,-2)

OpenStudy (cookiemonster2412):

@CausticSyndicalist @ilovewolf @poopsiedoodle

OpenStudy (cookiemonster2412):

@LiteNing1337

OpenStudy (litening1337):

Oh damn, I suck at Math. XD

OpenStudy (cookiemonster2412):

xD

OpenStudy (ilovewolf):

since you wanna be tagged @hysusonic

OpenStudy (anonymous):

Can you calculate the slope of the line from the two points given?

OpenStudy (cookiemonster2412):

@TheRaggedyDoctor

OpenStudy (cookiemonster2412):

THE SLOPE IS -1/3

OpenStudy (hysusonic):

first you plug in the numbers for y=mx+b (the equation of the line) so you already have the m, now you need the b

OpenStudy (cookiemonster2412):

and how do i do that????

OpenStudy (hysusonic):

m=slope b=y-intercept

OpenStudy (litening1337):

That means where on the Y-axis is the line going through?

OpenStudy (cookiemonster2412):

i have to use the point slope formula y-y1=m(x-x1)

OpenStudy (anonymous):

You are given \[\left( x_1,y_1 \right) \text{ and }\left( x_2,y_2) \right)\]Calculate slope m using\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]Then substitute m, y1 and x1 into the point-sloep form.

OpenStudy (cookiemonster2412):

y-(-1)=-1/3(x-(-3) thats all i have

OpenStudy (anonymous):

Very good. You can still simplify a bit. Try to simplify y--1 and x--3.

OpenStudy (cookiemonster2412):

y=-1/3x-2?

OpenStudy (anonymous):

When you subtract a negative number, it becomes an addition. For example, x-(-5) simplifies to x+5. There are two instances of this in your earlier solution. Try simplifying them as I've shown here.

OpenStudy (cookiemonster2412):

so im wrong

OpenStudy (anonymous):

Yes, y=-1/3x-2 is incorrect. Back up one step and simplify where you are subtracting negative numbers